OpenStudy (anonymous):

Convert 24 cm to the unit km. 0.00024 km 0.24 km 2.4 x 105 km 2.4 x 10-3 km @photon336

OpenStudy (photon336):

We convert centimeters to meters by using the conversion factor below: \[24 cm *(\frac{ meter }{ 100cm }) = \frac{ 24 }{ 100 } = 0.24meter\] then we convert meters to kilometers by multiplying meters by the conversion factor below. notice that the units for meters cancel out and we're left with kilometers as the answer. \[0.24meters*(\frac{ Km }{ 1000meters }) = (\frac{ 2.4x10^{-1} }{ 1.0x10^{3}}) = 2.4x10^{-4} km \]

OpenStudy (photon336):

\[2.4x10^{-4} = 0.00024\]

OpenStudy (anonymous):

What is the main energy conversion that occurs when you plug in a lamp? Electrical energy is converted to nuclear energy. Thermal energy is converted to electrical energy. Electrical energy is converted to radiant energy. Nuclear energy is converted to radiant energy. @photon336

OpenStudy (anonymous):

A) 0.00024km

OpenStudy (anonymous):

C) Electrical energy is converted to radiant energy.